BRICS Research Series, 1998

August 5, 2003

This document is also available as PostScript and DVI.

References

[RS-98-55] Abstract, PostScript, PDF, DVI.

Andrew D. Gordon, Paul D. Hankin, and Sren B. Lassen. Compilation and Equivalence of Imperative Objects (Revised Report). December 1998. iv+75 pp. This is a revision of Technical Report 429, University of Cambridge Computer Laboratory, June 1997, and the earlier BRICS report RS-97-19, July 1997. Appears in Journal of Functional Programming, 9(4):373-427, 1999, and Ramesh and Sivakumar, editors, Foundations of Software Technology and Theoretical Computer Science: 17th Conference, FST&TCS ’97 Proceedings, LNCS 1346, 1997, pages 74–87.

[RS-98-54] Olivier Danvy and Ulrik P. Schultz. Lambda-Dropping: Transforming Recursive Equations into Programs with Block Structure. December 1998. 55 pp. This report is superseded by the later report BRICS RS-99-27.

[RS-98-53] Abstract, PostScript, PDF.

Julian C. Bradfield. Fixpoint Alternation: Arithmetic, Transition Systems, and the Binary Tree. December 1998. 20 pp. Appears in Theoretical Informatics and Applications, 33:341–356, 1999.

[RS-98-52] Abstract, PostScript, PDF, DVI.

Josva Kleist and Davide Sangiorgi. Imperative Objects and Mobile Processes. December 1998. 22 pp. Appears in Gries and de Roever, editors, IFIP Working Conference on Programming Concepts and Methods, PROCOMET ’98 Proceedings, 1998, pages 285–303.

[RS-98-51] Abstract, PostScript, PDF.

Peter Krogsgaard Jensen. Automated Modeling of Real-Time Implementation. December 1998. 9 pp. Appears in The 13th IEEE Conference on Automated Software Engineering, ASE ’98 Doctoral Symposium Proceedings, 1998, pages 17–20.

[RS-98-50] Abstract, PostScript, PDF, DVI.

Luca Aceto and Anna Inglfsdttir. Testing Hennessy-Milner Logic with Recursion. December 1998. 15 pp. Appears in Thomas, editor, Foundations of Software Science and Computation Structures: Second International Conference, FoSSaCS ’99 Proceedings, LNCS 1578, 1999, pages 41–55.

[RS-98-49] Abstract, PostScript, PDF, DVI. Luca Aceto, Willem Jan Fokkink, and Anna Inglfsdttir. A Cook’s Tour of Equational Axiomatizations for Prefix Iteration. December 1998. 14 pp. Appears in Nivat, editor, Foundations of Software Science and Computation Structures: First International Conference, FoSSaCS ’98 Proceedings, LNCS 1378, 1998, pages 20–34.

[RS-98-48] Abstract, PostScript, PDF, DVI. Luca Aceto, Patricia Bouyer, Augusto Burgueo, and Kim G. Larsen. The Power of Reachability Testing for Timed Automata. December 1998. 12 pp. Appears in Arvind and Ramanujam, editors, Foundations of Software Technology and Theoretical Computer Science: 18th Conference, FST&TCS ’98 Proceedings, LNCS 1530, 1998, pages 245–256.

[RS-98-47] Abstract, PostScript, PDF. Gerd Behrmann, Kim G. Larsen, Justin Pearson, Carsten Weise, and Wang Yi. Efficient Timed Reachability Analysis using Clock Difference Diagrams. December 1998. 13 pp. Appears in Halbwachs and Peled, editors, Computer-Aided Verification: 11th International Conference, CAV ’99 Proceedings, LNCS 1633, 1999, pages 341–353.

[RS-98-46] Abstract, PostScript, PDF. Kim G. Larsen, Carsten Weise, Wang Yi, and Justin Pearson. Clock Difference Diagrams. December 1998. 18 pp. Presented at 10th Nordic Workshop on Programming Theory, NWPT ’10 Abstracts, Turku Centre for Computer Science TUCS General Publications 11, 1998. Appears in Nordic Journal of Computing, 6(3):271–298, 1999.

[RS-98-45] Abstract, PostScript, PDF. Morten Vadskr Jensen and Brian Nielsen. Real-Time Layered Video Compression using SIMD Computation. December 1998. 37 pp. Appears in Zinterhof, Vajtersic and Uhl, editors, Parallel Computing: Fourth International ACPC Conference, ACPC ’99 Proceedings, LNCS 1557, 1999, pages 377–387.

[RS-98-44] Abstract, PostScript, PDF. Brian Nielsen and Gul Agha. Towards Re-usable Real-Time Objects. December 1998. 36 pp. Appears in The Annals of Software Engineering, 7:257–282, 1999.

[RS-98-43] Abstract, PostScript, PDF, DVI. Peter D. Mosses. CASL: A Guided Tour of its Design. December 1998. 31 pp. Appears in Fiadeiro, editor, Recent Trends in Algebraic Development Techniques: 13th Workshop, WADT ’98 Selected Papers, LNCS 1589, 1999, pages 216–240.

[RS-98-42] Abstract, PostScript, PDF, DVI. Peter D. Mosses. Semantics, Modularity, and Rewriting Logic. December 1998. 20 pp. Appears in Kirchner and Kirchner, editors, International Workshop on Rewriting Logic and its Applications, WRLA ’98 Proceedings, ENTCS 15, 1998.

[RS-98-41] Abstract, PostScript, PDF, DVI. Ulrich Kohlenbach. The Computational Strength of Extensions of Weak Knig’s Lemma. December 1998. 23 pp.

[RS-98-40] Abstract, PostScript, PDF, DVI. Henrik Reif Andersen, Colin Stirling, and Glynn Winskel. A Compositional Proof System for the Modal μ-Calculus. December 1998. 29 pp. Extended abstract appears in Ninth Annual IEEE Symposium on Logic in Computer Science, LICS ’94 Proceedings, 1994, pages 144–153. Superseeds the earlier BRICS Report RS-94-34.

[RS-98-39] Abstract, PostScript, PDF, DVI. Daniel Fridlender. An Interpretation of the Fan Theorem in Type Theory. December 1998. 15 pp. Appears in Altenkirch, Naraschewski and Reus, editors, International Workshop on Types for Proofs and Programs 1998, TYPES ’98 Selected Papers, LNCS 1657, 1999, pages 93–105.

[RS-98-38] Abstract, PostScript, PDF, DVI. Daniel Fridlender and Mia Indrika. An n-ary zipWith in Haskell. December 1998. 12 pp.

[RS-98-37] Abstract, PostScript, PDF, DVI. Ivan B. Damgrd, Joe Kilian, and Louis Salvail. On the (Im)possibility of Basing Oblivious Transfer and Bit Commitment on Weakened Security Assumptions. December 1998. 22 pp. Appears in Stern, editor, Advances in Cryptology: International Conference on the Theory and Application of Cryptographic Techniques, EUROCRYPT ’99 Proceedings, LNCS 1592, 1999, pages 56–73.

[RS-98-36] Abstract, PostScript, PDF, DVI. Ronald Cramer, Ivan B. Damgrd, Stefan Dziembowski, Martin Hirt, and Tal Rabin. Efficient Multiparty Computations with Dishonest Minority. December 1998. 19 pp. Appears in Stern, editor, Advances in Cryptology: International Conference on the Theory and Application of Cryptographic Techniques, EUROCRYPT ’99 Proceedings, LNCS 1592, 1999, pages 311–326.

[RS-98-35] Abstract.

Olivier Danvy and Zhe Yang. An Operational Investigation of the CPS Hierarchy. December 1998. Extended version of a paper appearing in Swierstra, editor, Programming Languages and Systems: Eighth European Symposium on Programming, ESOP ’99 Proceedings, LNCS 1576, 1999, pages 224–242.

[RS-98-34] Abstract, PostScript, PDF, DVI. Peter G. Binderup, Gudmund Skovbjerg Frandsen, Peter Bro Miltersen, and Sven Skyum. The Complexity of Identifying Large Equivalence Classes. December 1998. Appears in Fundamenta Informaticae, 38:17–29.

[RS-98-33] Abstract, PostScript, PDF, DVI. Hans Httel, Josva Kleist, Uwe Nestmann, and Massimo Merro. Migration = Cloning ; Aliasing (Preliminary Version). December 1998. 40 pp. Appears in Cardelli, editor, Foundations of Object-Oriented Languages: 6th International Conference, FOOL6 Informal Proceedings, 1999.

[RS-98-32] Abstract, PostScript, PDF, DVI. Jan Camenisch and Ivan B. Damgrd. Verifiable Encryption and Applications to Group Signatures and Signature Sharing. December 1998. 18 pp. Appears in Okamoto, editor, Advances in Cryptology: Sixth ASIACRYPT Conference on the Theory and Applications of Cryptologic Techniques, ASIACRYPT ’00 Proceedings, LNCS 1976, 2000, pages 331–345.

[RS-98-31] Abstract, PostScript, PDF, DVI. Glynn Winskel. A Linear Metalanguage for Concurrency. November 1998. 21 pp. Appears in Haeberer, editor, Algebraic Methodology and Software Technology: 7th International Conference, AMAST ’98 Proceedings, LNCS 1548, 1999, pages 42–58.

[RS-98-30] Abstract, PostScript, PDF, DVI. Carsten Butz. Finitely Presented Heyting Algebras. November 1998. 30 pp.

[RS-98-29] Abstract, PostScript, PDF, DVI. Jan Camenisch and Markus Michels. Proving in Zero-Knowledge that a Number is the Product of Two Safe Primes. November 1998. 19 pp. Appears in Stern, editor, Advances in Cryptology: International Conference on the Theory and Application of Cryptographic Techniques, EUROCRYPT ’99 Proceedings, LNCS 1592, 1999, pages 106–121.

[RS-98-28] Abstract, PostScript, PDF. Rasmus Pagh. Low Redundancy in Dictionaries with O(1) Worst Case Lookup Time. November 1998. 15 pp. Journal version in SIAM Journal on Computing 31:353–363, 2001, and proceedings version in van Emde Boas, Wiedermann and Nielsen, editors, 26th International Colloquium on Automata, Languages, and Programming, ICALP ’99 Proceedings, LNCS 1644, 1999, pages 595–604.

[RS-98-27] Abstract, PostScript, PDF, DVI. Jan Camenisch and Markus Michels. A Group Signature Scheme Based on an RSA-Variant. November 1998. 18 pp. Preliminary version appeared in Ohta and Pei, editors, Advances in Cryptology: Fourth ASIACRYPT Conference on the Theory and Applications of Cryptologic Techniques, ASIACRYPT ’98 Proceedings, LNCS 1514, 1998, pages 160–174.

[RS-98-26] Abstract, PostScript, PDF, DVI. Paola Quaglia and David Walker. On Encoding pπ in mπ. October 1998. 27 pp. Full version of paper appearing in Arvind and Ramanujam, editors, Foundations of Software Technology and Theoretical Computer Science: 18th Conference, FST&TCS ’98 Proceedings, LNCS 1530, 1998, pages 42–53.

[RS-98-25] Abstract, PostScript, PDF, DVI. Devdatt P. Dubhashi. Talagrand’s Inequality in Hereditary Settings. October 1998. 22 pp.

[RS-98-24] Abstract, PostScript, PDF, DVI. Devdatt P. Dubhashi. Talagrand’s Inequality and Locality in Distributed Computing. October 1998. 14 pp Appears in Luby, Rolim and Serna, editors, Randomization and Approximation Techniques in Computer Science: Second International Workshop: Second International Workshop, RANDOM ’98 Proceedings, LNCS 1518, 1998, pages 60–70.

[RS-98-23] Abstract, PostScript, PDF, DVI. Devdatt P. Dubhashi. Martingales and Locality in Distributed Computing. October 1998. 19 pp. Appears in Arvind and Ramanujam, editors, Foundations of Software Technology and Theoretical Computer Science: 18th Conference, FST&TCS ’98 Proceedings, LNCS 1530, 1998, pages 174–185.

[RS-98-22] Abstract, PostScript, PDF, DVI. Gian Luca Cattani, John Power, and Glynn Winskel. A Categorical Axiomatics for Bisimulation. September 1998. ii+21 pp. Appears in Sangiorgi and de Simone, editors, Concurrency Theory: 9th International Conference, CONCUR ’98 Proceedings, LNCS 1466, 1998, pages 581–596.

[RS-98-21] Abstract, PostScript, PDF, DVI. John Power, Gian Luca Cattani, and Glynn Winskel. A Representation Result for Free Cocompletions. September 1998. 16 pp. Appears in Journal of Pure and Applied Algebra, 151(3):273–286 (2000).

[RS-98-20] Abstract, PostScript, PDF, DVI. Sren Riis and Meera Sitharam. Uniformly Generated Submodules of Permutation Modules. September 1998. 35 pp. Appears in Annals of Pure and Applied Algebra, 160(3):285–318, 2001.

[RS-98-19] Abstract, PostScript, PDF, DVI. Sren Riis and Meera Sitharam. Generating Hard Tautologies Using Predicate Logic and the Symmetric Group. September 1998. 13 pp. Abstract appears in Logic Journal of IGPL, a special section devoted the 5th Workshop on Logic, Language, Information and Computation, WoLLIC ’98 (So Paulo, Brazil, July 28–31, 1998), pages 936– 937. Full version appears in Logic Journal of the IGPL, 8(6):787–795, 2000.

[RS-98-18] Abstract, PostScript, PDF, DVI.

Ulrich Kohlenbach. Things that can and things that can’t be done in PRA. September 1998. 24 pp. Appeared in Annals of Pure and Applied Logic, 102(3):223-245, 2000.

[RS-98-17] Abstract, PostScript, PDF, DVI.

Roberto Bruni, Jos Meseguer, Ugo Montanari, and Vladimiro Sassone. A Comparison of Petri Net Semantics under the Collective Token Philosophy. September 1998. 20 pp. Appears in Hsiang and Ohori, editors, Fourth Asian Computing Science Conference, ASIAN ’98 Proceedings, LNCS 1538, 1998, pages 225–244.

[RS-98-16] Stephen Alstrup, Thore Husfeldt, and Theis Rauhe. Marked Ancestor Problems. September 1998. Extended abstract appears in Motwani, editor, 39th Annual Symposium on Foundations of Computer Science, FOCS ’98 Proceedings, 1998, pages 534–543.

[RS-98-15] Abstract, PostScript, PDF.

Jung-taek Kim, Kwangkeun Yi, and Olivier Danvy. Assessing the Overhead of ML Exceptions by Selective CPS Transformation. September 1998. 31 pp. Appears in the proceedings of the 1998 ACM SIGPLAN Workshop on ML, Baltimore, Maryland, September 26, 1998, pages 103–114.

[RS-98-14] Abstract, PostScript, PDF, DVI.

Sandeep Sen. The Hardness of Speeding-up Knapsack. August 1998. 6 pp.

[RS-98-13] Abstract, PostScript, PDF, DVI.

Olivier Danvy and Morten Rhiger. Compiling Actions by Partial Evaluation, Revisited. June 1998. 25 pp. Earlier version appears in Vain, editor, 9th Nordic Workshop on Programming Theory, NWPT ’9 Proceedings, Estonian Academy of Sciences Series, 1997.

[RS-98-12] Abstract, PostScript, PDF, DVI.

Olivier Danvy. Functional Unparsing. May 1998. 7 pp. This report supersedes the earlier report BRICS RS-98-5. Extended version of an article appearing in Journal of Functional Programming, 8(6):621– 625, 1998.

[RS-98-11] Abstract, PostScript, PDF, DVI.

Gudmund Skovbjerg Frandsen, Johan P. Hansen, and Peter Bro Miltersen. Lower Bounds for Dynamic Algebraic Problems. May 1998. 30 pp. Revised versions appears in Meinel and Tison, editors, 16th Annual Symposium on Theoretical Aspects of Computer Science Proceedings, STACS ’99 Proceedings, LNCS 1563, 1999, pages 362– 372, and in Information and Computation, 171:333–349, 2001.

[RS-98-10] Abstract, PostScript, PDF, DVI.

Jakob Pagter and Theis Rauhe. Optimal Time-Space Trade-Offs for Sorting. May 1998. 12 pp. Appears in Motwani, editor, 39th Annual Symposium on Foundations of Computer Science, FOCS ’98 Proceedings, 1998, pages 264–268.

[RS-98-9] Abstract, PostScript, PDF.

Zhe Yang. Encoding Types in ML-like Languages (Preliminary Version). April 1998. 32 pp. Extended version of a paper appearing in Hudak and Queinnec, editors, Proceedings of the third ACM SIG-PLAN International Conference on Functional Programming, 1998, pages 289–300.

[RS-98-8] Abstract, PostScript, PDF, DVI.

P. S. Thiagarajan and Jesper G. Henriksen. Distributed Versions of Linear Time Temporal Logic: A Trace Perspective. April 1998. 49 pp. Appears as a chapter of Reisig and Rozenberg, editors, Lectures on Petri Nets I: Basic Models, LNCS 1491, 1998, pages 643– 681.

[RS-98-7] Abstract, PostScript, PDF, DVI.

Stephen Alstrup, Thore Husfeldt, and Theis Rauhe. Marked Ancestor Problems (Preliminary Version). April 1998. 36 pp. Superseeded by RS-98-16. Extended abstract in Motwani, editor, 39th Annual Symposium on Foundations of Computer Science, FOCS ’98 Proceedings, 1998, pages 534–543.

[RS-98-6] Abstract, PostScript, PDF, DVI.

Kim Sunesen. Further Results on Partial Order Equivalences on Infinite Systems. March 1998. 48 pp.

[RS-98-5] Abstract, PostScript, PDF, DVI.

Olivier Danvy. Formatting Strings in ML. March 1998. 3 pp. This report is superseded by the later report BRICS RS-98-12.

[RS-98-4] Abstract, PostScript, PDF.

Mogens Nielsen and Thomas S. Hune. Timed Bisimulation and Open Maps. February 1998. 27 pp. Appears in Brim, Gruska and Zlatuška, editors, Mathematical Foundations of Computer Science: 23rd International Symposium, MFCS ’98 Proceedings, LNCS 1450, 1998, pages 378–387.

[RS-98-3] Abstract, PostScript, PDF, DVI.

Christian N. S. Pedersen, Rune B. Lyngs, and Jotun Hein. Comparison of Coding DNA. January 1998. 20 pp. Appears in Farach-Colton, editor, Combinatorial Pattern Matching: 9th Annual Symposium, CPM ’98 Proceedings, LNCS 1448, 1998, pages 153–173.

[RS-98-2] Olivier Danvy. An Extensional Characterization of Lambda-Lifting and Lambda-Dropping. January 1998. This report is superseded by the later report BRICS RS-99-21.

[RS-98-1] Abstract, PostScript, PDF, DVI.

Olivier Danvy. A Simple Solution to Type Specialization. January 1998. 10 pp. Appears in Larsen, Skyum and Winskel, editors, 25th International Colloquium on Automata, Languages, and Programming, ICALP ’98 Proceedings, LNCS 1443, 1998, pages 908–917.